OpenStudy (anonymous):

Calculate the concentration of potassium ion in grams per liter after mixing 100ml of 0.250M KCl and 200ml of 0.1 M K2SO4

OpenStudy (anonymous):

first you write dissociation of KCl and K2SO4 where you assume that KCl and K2SO4 are completely dissasociated, then you set the ratio for each K+ ion and its compound and from there get the amount of K+ ions from each compound, sum them up and calculate mass, after that divide with summed volume and there you go.

OpenStudy (anonymous):

What exactly do I sum up?

OpenStudy (anonymous):

KCl -> K+ + Cl- K2SO4 -> 2K+ + SO4 2- c= n/V -> n= c*V n(KCl) = c(KCl) * V(KCl) n1(K+)/n(KCl) = 1/1 = 1 n1(K+) = n(KCl) n(K2SO4) = c(K2SO4) * V(K2SO4) n2(K+)/n(K2SO4) = 2/1 = 2 n2(K+) = 2*n(K2SO4) n(K+) = n1(K+) + n2(K+) n= m/M -> m = n*M m(K+) = n(K+) * M(K) V = V(KCl) + V(K2SO4) gama - mass concentration gama = m(K+) / V

OpenStudy (anonymous):

thank you!